1. You are working in the Emergency Room and haloperidol 7.5 mg IM is ordered in agitated patients with schizophrenia.

Label reads;

Haloperidol 50mg/10mL

How many mL will you give?_______mL (answer to the nearest tenth of an mL)

2. The physician has orders for a patient with generalized anxiety disorder lorazepam (Ativan) 0.02 mg/kg every 3 hours IV PRN. The patient weight is 55kg. Lorazepam is available as 2mg/mL. How many mg do you give for each dose? ____mg. (Round to the nearest 10th)

3. There are essential aspects of care for a patient who is experiencing a tonic-clonic seizure? Which answer is incorrect?

a. Help the person to the floor.

b. Loosen clothing around the neck

c. Turn the person to the side

d. Remove any sharp objects

e. Administer an immediate oral dose of antiseizure medication.
Answer #1

1. ordered haloperidol = 7.5 mg IM

Labelled Haloperidol= 50mg/10mL

menas 5 mg/ml

for 7.5 mg required ml= 7.5/5= 1.5 ml should be given

2.. Ortdered  lorazepam (Ativan) = 0.02 mg/kg every 3 hours IV PRN.

The patient weight is 55kg.

Required dose as per weight of the patient will be = 55X 0.02= 1.1 mg

Available Lorazepam= 2mg/mL.

Required ml per dose will be= 1.1/2 ml= 0.52 ml

3. Except option E all are ther essential aspects of care for a patient who is experiencing a tonic-clonic seizure like loosening of the cloth, helping the person to the floor, removing the shap object as it be the cause of injury as well as turn the person on side